Reports 1-1 of 1 Clear search Modify search
takafumi.ushiba - 23:34 Wednesday 10 August 2022 (21753) Print this report
PR3 OpLev rearangement

[Yokozawa, Yasui, Ushiba]

We rearanged PR3 OpLevs to avoid sticking.out QPDs form Optical tables on the pylon.
Rearangement was succeeded and new wind shield made by aluminum foils were constracted.

Since PR3 OpLev output beam was passing through the very edge of the viewport and QPDs for the tilt OpLev was installed sticking out from the optical table.
So, we moved injection beam position and rearaged optical layout to avoid protrusion.

<What we did>
We moved injection beam position and adjusted the injection angle in order not to clip the beam.
After that, we rearange the reciever side optical components.
To avoid changing calibration factors, we set a lens, mirror, and QPD for length sensing OpLev as follows:
*Distance between BS and lens: 25 mm
*Distance between lens and mirror: 215 mm
*Distance between mirror and QPD: 215 mm
According to the klog19613, total diatance from BS to QPD is 455 mm, which is the same as the current setup.

After that, we constracted wind shield for OpLevs.
In this time, we didn't make windshield between viewport flanges and pylons because we will soon have leak test of central area after vacuum evacuation.
So, we need to make it after cnfirming vacuum leak was not there.
In addition, we made aluminum foil wind shield so as to be easily open by only detaching white tapes.
So, please do not touch pink tapes on the windshield, otherwise the windshield will be broken and need to be repaired from the begining. 

After the work, I recovered the alignment and current setpoint of PR3 is a good enough to lock the X arm.

Comments to this report:
takafumi.ushiba - 23:41 Wednesday 10 August 2022 (21754) Print this report

[Yamamoto, Ushiba]

After this work, we found one of the IM OSEM signal cable was disconnected and cannot reach LOCK_ACQUISITION state.
This situation was already recovered but please do not leave the screws of cables loose, or this kind of happening won't go away from KAGRA.

At least on PR3 satellite box, we found three unlocked cables and five loose cables out of 15; it's terrible.

takaaki.yokozawa - 17:40 Thursday 11 August 2022 (21755) Print this report
Some additional comments.

1. The screw of length sensing QPD between circuit and QPD stand was too loose (Maybe tightened the screw just by hand), it made the unstable setting of the QPD, be careful.
2. Rotate the BS 45 degree to minimize the effect of ghost beam from mirror wedge
3. We added the IR, GR filter and cover to TILT QPD.
- When we set the cover to TILT QPD, we noticed we couldn't move QPD in vertical axis, touching the cover to micromater
- Ushiba-san found that it may happened the relative position in the QPD stand due to the small clearance in the screw, after small adjustment of the QPD stand, its problem was solved
- We couldn't set the cover to length sensing QPD, so we need to discuss among the oplev experts
4. I placed some pictures
5. (This is just a homework) we need to check the sensitivity of length sensing QPD by oscillating the PR3 suspension by pitch motion
6. (This is just a homework) we will check the performance of IR/GR filter of the angle dependence
Images attached to this comment
Search Help